OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 625 and a3 = 25?

OpenStudy (anonymous):

is a1 the 1st term and a3 the 3rd term?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

tn=ar^n-1 t3=(625)(r^3-1) 25=625r^2 1/25 = r^2 1/5=r tn=ar^n-1 t6= (625)(1/5 ^5) t6=625x 1/3125 t6=0.2